# version 1.0

### v6

@@ 3.4 The Value Iteration Algorithm
In the previous section, we established the Bellman equation and its consequence to a polynomial-time solution for reinforcement learning in Markov Decision Processes. In this section, we study a fundamental algorithm for solving Markov Decision Processes that is a more direct application of dynamic programming and the Bellman equation.
The value iteration algorithm starts from an arbitrary solution and improves it until a certain measure of optimality is reached. A formal description is given in Algorithm 2. The performance guarantee of Value Iteration, as well as its analysis, is particularly elegant and does not require expertise in mathematical optimization theory.
@@
$\mathbf{v}_{t+1}(s)=\max_{a}\left\{R_{sa}+\gamma\sum_{s^{\prime}}P_{ss^{\prime}}^{a}\mathbf{v}_{t}(s^{\prime})\right\}$
@@
return $\pi_{T}$ which is derived from the last value function as
$\pi_{T}(s)=\operatorname*{\arg\max}_{a}\left\{R_{sa}+\gamma\sum_{s^{\prime}}P_{ss^{\prime}}^{a}\mathbf{v}_{T}(s^{\prime})\right\}$
Beyond being simple and providing a finite-time convergence rate, the analysis also establishes feasibility of the optimal solution, as given in the following theorem.
@@
Define the residual vector $\mathbf{r}_{t}(s)=\mathbf{v}^{*}(s)-\mathbf{v}_{t}(s)$. Then the optimal value vector $\mathbf{v}^{*}$ is unique and the value iteration converges to the optimal value function and policy at a geometric rate of
$\|\mathbf{r}_{t+1}\|_{\infty}\leq\gamma^{t}\|\mathbf{r}_{1}\|_{\infty}.$
@@ 3.5 Bibliographic Remarks
The section on reinforcement learning in the text [RN02] introduces basic concepts, along with commentaries on related topics, such as the origins of the scalar reward model from utility theory. The classical text on reinforcement learning, authored by pioneers of the subject, is [SB18]. In addition, this text extensively discusses early ventures into the field of researchers working in artificial intelligence, psychology, and neuroscience.
Another great reference is the set of lecture notes [Sil15]. See [AJKS19] for an excellent introduction to the theoretical aspects of reinforcement learning, such as sample complexity bounds and reductions to supervised learning. The reader may also refer to [Ber19] for an extensive survey of computational approaches to learning in unknown Markov decision processes.
The first author had the benefit of a discussion with [Aro20] on the proof of convergence of value iteration presented in this chapter.
Dynamic programming approaches, in addition to being widely used in algorithm design [CLRS22], are at the heart of Markov decision processes. This view is emphasized in [Ber07], which discusses the sufficiency and necessity of the Bellman optimality conditions at length. It is also a good resource for an involved reader who might be interested in the completeness of the Bellman characterization and the existential concerns for the average-time setting, which we did not expand upon.
Although tabular MDPs do not compactly represent stochastic linear dynamical systems, function-approximation schemes [JYWJ20, DKL+21, JKA+17, KKL+20] successfully capture the stochastic LQR setting.
@@ 3.6 Exercises
3.1. In this exercise, we give a guided proof of the ergodic theorem. Consider an irreducible and aperiodic Markov chain with a transition matrix $P$.
A. Show that $P$ has a left eigenvalue equal to 1.
B. Show that all eigenvalues of $P$ are at most one in magnitude.
The irreducibility and aperiodicity of the Markov chain imply that the eigenvalue equal to one is unique. We are going to assume this fact and denote by $\mathbf{v}_{1}$ the corresponding eigenvector.
C. Consider an abirtrary distribution $\mathbf{p}_{0}$, and the distribution after $t$ iterations of the Markov chain $\mathbf{p}_{t}=\mathbf{p}_{0}P^{t}$. Prove that $\lim_{t\to\infty}\mathbf{p}_{t}=\mathbf{v}_{1}.$
3.2. Prove the geometric convergence of the value iteration algorithm for rewards that can be arbitrary real numbers, as opposed to in the range $[0,1]$.
3.3. Consider the derivation of an optimal policy for a Markov Decision Process given its value function as described in algorithm 2. Prove that the resulting policy is indeed optimal.
3.4. Let Markov chain $P$ be reversible with respect to stationary distribution $\boldsymbol{\pi}$}, i.e.
@@
3.5. Prove that the successive iterates $\mathbf{v}_{t}$ produced by the Value Iteration algorithm increase monotonically, that is, $\mathbf{v}_{t+1}\geq\mathbf{v}_{t}$, when initialized with $\mathbf{v}_{0}=\mathbf{0}$. Hence, conclude that for any $t$, $\mathbf{v}^{*}\geq\mathbf{v}_{t}$.
3.6. Let $P$ be an irreducible and aperiodic Markov chain. Prove that $P$ is reversible if and only if for any sequence of states $s_{1},\dots s_{n}$, it holds that
$P_{s_{1}s_{2}}P_{s_{2}s_{3}}\dots P_{s_{n-1}s_{n}}P_{s_{n}s_{1}}=P_{s_{1}s_{n}}P_{s_{n}s_{n-1}}\dots P_{s_{2}s_{1}}.$
@@
In Chapter 2, we observed that even simple questions about general dynamical systems are intractable. This motivates the restriction to a special class of dynamical systems for which we can design efficient algorithms with provable guarantees.
@@
If the matrices ${A}_{t}={A}$ and ${B}_{t}={B}$ are fixed, we say that the dynamical system is linear time-invariant (LTI). We have seen examples of LTI systems in Chapter 2, namely the double integrator and the aircraft dynamics.
@@
A more general class of linear dynamical systems is when the matrices ${A}_{t}$ and ${B}_{t}$ change over time. These are called linear time-varying (LTV) systems. The class of LTV systems can express more general dynamics via linearization, as we show next.
@@ 4.1 Approximating General Dynamics via LTV Systems
Consider any dynamics that can be expressed by a smooth function $f$. Such a system can be approximated by a time-varying linear system to an arbitrary accuracy determined by the time-step interval between samples.
To see this, first consider the one dimensional first-order Taylor expansion of a given dynamics function $f$, which is twice differentiable, around an anchor point $(\bar{x}_{t},\bar{u}_{t})$:
$x_{t+1}$ $=f(x_{t},u_{t})+w_{t}$
$=f(\bar{x}_{t},\bar{u}_{t})+\underbrace{\frac{\partial}{\partial x}f(x,u)\bigg|_{\bar{x}_{t},\bar{u}_{t}}}_{\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}A_{t}}(x_{t}-\bar{x}_{t})+\underbrace{\frac{\partial}{\partial u}f(x,u)\bigg|_{\bar{x}_{t},\bar{u}_{t}}}_{\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}B_{t}}(u_{t}-\bar{u}_{t})+w_{t},$
where $w_{t}$ is the approximation error, or remainder, which by the mean value theorem behaves as $w_{t}\sim(x_{t}-\bar{x}_{t})^{2}+(u-\bar{u}_{t})^{2}$. Thus, this simplifies to
$x_{t+1}=A_{t}x_{t}+B_{t}u_{t}+w_{t}.$
An analogous approximation holds for the higher-dimensional case. The Taylor expansion is
$f(\mathbf{x},\mathbf{u})=f(\bar{\mathbf{x}},\bar{\mathbf{u}})+J_{\mathbf{x}}f(\bar{\mathbf{x}},\bar{\mathbf{u}})(\mathbf{x}-\bar{\mathbf{x}})+J_{\mathbf{u}}f(\bar{\mathbf{x}},\bar{\mathbf{u}})(\mathbf{u}-\bar{\mathbf{u}})+w_{t},$
where $J_{\mathbf{x}}f(\bar{\mathbf{x}},\bar{\mathbf{u}})$ is the Jacobian of $f$ with respect to variables $\bar{\mathbf{x}}$. Writing the matrices ${A}_{t},{B}_{t}$ for the Jacobians, we have
$\mathbf{x}_{t+1}={A}_{t}\mathbf{x}_{t}+{B}_{t}\mathbf{u}_{t}+\mathbf{w}_{t}.$
This derivation explains why in practice the class of time-varying linear dynamical systems is very expressive, at least for smooth dynamical systems. This is particularly true with hidden states that can capture complicated dynamics while imposing enough structure for efficient algorithms.
@@ 4.2 Stabilizability of Linear Systems
Verifying stability of a system, and more so stabilizability, was shown to be intractable in our discussion of chapter 2. For linear dynamical systems, however, the picture is quite different: stability and stabilizability can be efficiently verified! Furthermore, the theoretical analysis of this fact is interesting and elegant.
@@
Recall that a system is stabilizable if, from every starting state, there is a sequence of controls that drives the system to zero. Determining whether a linear time-invariant system is stabilizable is determined by a particularly simple condition as follows. Henceforth, we denote by $\rho({A})$ the spectral radius of a matrix A{A}, that is,
$\rho({A})=\max\{|\lambda_{1}|,...,|\lambda_{d}|\},$
the maximum modulus of all (possibly complex) eigenvalues. The stabilizability of a linear dynamical system can be characterized as follows.
@@
We therefore prove the sufficiency of this condition and leave the necessity as exercise 4.6. Both directions rely on the characterization of the spectral norm as follows:
@@
$\rho({A})<1$ if and only if $\lim_{t\mapsto\infty}{A}^{t}=0$. Furthermore, if $\rho({A})>1$, then for any norm, $\lim_{t\mapsto\infty}\|{A}^{t}\|=\infty$.
The proof of this fact is left as exercise 4.6. Based on this fact, we can now prove the sufficiency part of Theorem 4.1.
@@
Suppose that there exists a K{K} such that $\rho({A}+{B}{K})<1$. Then choose all controls to be $\mathbf{u}_{t}={K}\mathbf{x}_{t}$. The evolution of the system without perturbation is given by
$\mathbf{x}_{t}=({A}+{B}{K})\mathbf{x}_{t-1}=...=({A}+{B}{K})^{t}\mathbf{x}_{0}.$
Thus, using the previous fact concerning the spectral radius, we have
$\lim_{t\to\infty}\|\mathbf{x}_{t}\|=\lim_{t\to\infty}\|({A}+{B}{K})^{t}x_{0}\|=\|0\mathbf{x}_{0}\|=0.$
@@ 4.2.1 Efficient Verification of Stabilizability
The spectral radius of a matrix can be computed to an arbitrary accuracy in polynomial time and can be approximated by the power method. This implies that the stability of a given system can be efficiently verified.
To efficiently check if a given system is stabilizable, we make note of the following result due to Lyapunov. We use the notation $A\succeq 0$ to denote that the matrix $A$ is positive semidefinite, meaning that it is symmetric and all its eigenvalues are nonnegative.
@@
$\rho({A})<1$ holds if and only if there exists a matrix ${P}\in{\mathbb{R}}^{d_{x}\times d_{x}}$ such that ${P}\succeq{I}$ and ${P}\succ{A}^{\top}{P}{A}$.
@@
Say there exists a matrix $P$ such that $P\succeq I$ and $P\succ A^{\top}PA$, then to show that $\rho(A)<1$, it suffices to prove that $\lim_{t\to\infty}A^{t}=0$. To observe the latter, first note that there exists a positive constant $c<1$ such that $cP\succeq A^{\top}PA$, implying that $c^{t}P\succeq(A^{t})^{\top}PA^{t}\succeq(A^{t})^{\top}A^{t}$ via repeated applications of the previous inequality. Therefore, we have $\lim_{t\to\infty}(A^{t})^{\top}A^{t}=0$, and consequently $\lim_{t\to\infty}A^{t}=0$ since $\|A^{\top}\|\|A\|=\|A^{\top}A\|$.
For the other direction, given some matrix $A$ with $\rho(A)<1$, consider $P=\sum_{t=0}^{\infty}(A^{t})^{\top}A^{t}$. Since $I+A^{\top}PA, PP$ satisfies the required conclusions. The fact that $P$ as defined exists, that is, that it is finite, as long as $\rho(A)<1$ can be verified via the Jordan normal form of $A$, which in particular asserts that $\|A^{t}\|\leq O(t^{d_{x}}\rho(A)^{t})$. ∎
Using this observation, one can write a semidefinite mathematical program that checks if a given system is stabilizable, and if so produces a matrix $K$ that guarantees a spectral radius smaller than one.
Theorem 4.4.
There exists a matrix K{K} such that $\rho({A}+{B}{K})<1$ if and only if there exist matrices ${P}\in\mathcal{S}^{d_{x}\times d_{x}}$ and ${L}\in{\mathbb{R}}^{d_{u}\times d_{x}}$ such that
$\begin{bmatrix}{P}&({A}{P}+{B}{L})^{\top}\\ {A}{P}+{B}{L}&{P}\end{bmatrix}\succ 0,\qquad{P}\succeq{I}.$ (4.1)
Furthermore, given such matrices P{P} and L{L} satisfying Equation 4.1, the matrix ${K}^{\prime}={L}{P}^{-1}$ satisfies $\rho({A}+{B}{K}^{\prime})<1$.
Proof.
We begin by observing the following result, using the Schur complement, regarding the semi-definiteness of block matrices.
Lemma 4.5.
Define a block matrix M{M} as
${M}=\begin{bmatrix}{A}&{B}\\ {B}^{\top}&{C}\end{bmatrix}$. (4.2)
Then, ${M}\succ{0}$ if and only if ${A}\succ{0}$ and ${C}-{B}^{\top}{A}^{-1}{B}\succ{0}$.
Note that since the spectral radius purely involves the eigenvalues of a matrix, it is invariant under transposition, and we have $\rho((A+BK)^{\top})=\rho(A+BK)$. Now using this fact and Theorem 4.3, we know that the stabilizability of a linear system is equivalent to the existence of matrices K{K} and P{P} such that ${P}\succeq{I}$ and ${P}\succ{(A+BK)P(A+BK)^{\top}}$. The above lemma allows us to reformulate this cubic constraint to a quadratic constraint. The stabilizability of a system is equivalent to existence of matrices K{K} and P{P} such that
$\begin{bmatrix}{P}&{((A+BK)P)^{\top}}\\ {(A+BK)P}&{P}\end{bmatrix}\succ{0},\qquad{P}\succeq{I}.$
Lastly, we reparameterize this constraint as a linear semidefinite constraint as demonstrated below, by choosing L=KP{L}={KP}.
Notice that this is indeed a reformulation of the mathematical feasibility program, and not a relaxation, since the reparamertization ${K}={LP^{-1}}$ is invertible. ∎

### v8

@@ 3.4 The Value Iteration Algorithm
The Bellman optimality equation also leads to one of the most fundamental algorithms for discounted MDPs: value iteration. Starting from an initial estimate of the value function, the algorithm repeatedly applies the Bellman optimality update.
@@
$\mathbf{v}_{t+1}(s)=\max_{a\in A}\left\{R_{sa}+\gamma\sum_{s^{\prime}\in S}P_{ss^{\prime}}^{a}\,\mathbf{v}_{t}(s^{\prime})\right\}$
@@
 Define a greedy policy with respect to $\mathbf{v}_{T}$ by
$\pi_{T}(s)\in\arg\max_{a\in A}\left\{R_{sa}+\gamma\sum_{s^{\prime}\in S}P_{ss^{\prime}}^{a}\,\mathbf{v}_{T}(s^{\prime})\right\}$
return $\mathbf{v}_{T}$ and $\pi_{T}$
The next theorem gives the basic convergence guarantee.
@@
Let $(\mathbf{v}_{t})_{t\geq 0}$ be the sequence produced by Value Iteration with initialization $\mathbf{v}_{0}=\mathbf{0}$. Then:
  1. 1.
    $\mathbf{0}\leq\mathbf{v}_{t}\leq\mathbf{v}_{t+1}\leq\mathbf{v}^{*}$ entry-wise for every $t\geq 0$.
  2. 2.
    If we define the residual vector $\mathbf{r}_{t}=\mathbf{v}^{*}-\mathbf{v}_{t}$, then
    $\|\mathbf{r}_{t+1}\|_{\infty}\leq\gamma\|\mathbf{r}_{t}\|_{\infty}\qquad\text{for every }t\geq 0.$
    Consequently,
    $\|\mathbf{v}^{*}-\mathbf{v}_{t}\|_{\infty}\leq\frac{\gamma^{t}}{1-\gamma}.$
In particular, $\mathbf{v}_{t}\to\mathbf{v}^{*}$ geometrically fast.
@@
Monotonicity and the upper bound are left as an exercise. We now prove the geometric decay of the residual.
@@
Finally, because rewards lie in $[0,1]$, the total discounted reward is at most $1+\gamma+\gamma^{2}+\cdots=1/(1-\gamma)$, so
$\|\mathbf{r}_{0}\|_{\infty}=\|\mathbf{v}^{*}\|_{\infty}\leq\frac{1}{1-\gamma}.$
Combining the two inequalities yields
$\|\mathbf{v}^{*}-\mathbf{v}_{t}\|_{\infty}=\|\mathbf{r}_{t}\|_{\infty}\leq\frac{\gamma^{t}}{1-\gamma}.$
@@ 3.5 Bibliographic Remarks
The section on reinforcement learning in the text [RN02] introduces basic concepts and places them in the broader context of artificial intelligence. The classical textbook on reinforcement learning is [SB18], which develops the subject from both the algorithmic and conceptual perspectives.
Another excellent reference is the set of lecture notes [SIL15]. See [AJK+19a] for an introduction to the theoretical aspects of reinforcement learning, including sample-complexity guarantees and reductions to supervised learning. The reader may also refer to [BER19] for a broad survey of computational approaches to learning in unknown Markov decision processes.
Dynamic programming is at the heart of Markov decision processes. This viewpoint is emphasized in [BER07], which discusses the Bellman equations and related algorithmic ideas at length. It is also a useful source for the finite-horizon and average-reward settings, which we have only mentioned briefly here.
Although tabular MDPs do not compactly represent stochastic linear dynamical systems, function-approximation schemes [JYW+20, DKL+21, JKA+17, KKL+20] can successfully capture problems such as stochastic LQR.
@@ 3.6 Exercises
3.1. In this exercise, we give a guided proof of the convergence part of the ergodic theorem. Consider a finite irreducible and aperiodic Markov chain with transition matrix $P$.
A. Show that the all-ones vector is a right eigenvector of $P$ with eigenvalue $1$. Conclude that $1$ is also an eigenvalue of $P^{\top}$.
B. Show that all eigenvalues of $P$ have magnitude at most one.
The irreducibility and aperiodicity of the Markov chain imply that the eigenvalue $1$ of $P^{\top}$ is simple and that all other eigenvalues have magnitude strictly smaller than one. Assume this fact, and let $\mathbf{v}_{1}$ denote the corresponding left eigenvector normalized to sum to one.
C. Consider an arbitrary initial distribution $\mathbf{p}_{0}$, and the distribution after $t$ iterations of the Markov chain, $\mathbf{p}_{t}=\mathbf{p}_{0}P^{t}$. Prove that
$\lim_{t\to\infty}\mathbf{p}_{t}=\mathbf{v}_{1}.$
3.2. Prove the geometric convergence of the value-iteration algorithm for rewards that can be arbitrary bounded real numbers, as opposed to lying in the range $[0,1]$.
3.3. Let $\mathbf{v}^{*}$ be the optimal value function of a discounted Markov decision process, and define a greedy policy by
$\pi(s)\in\arg\max_{a\in A}\left\{R_{sa}+\gamma\sum_{s^{\prime}\in S}P^{a}_{ss^{\prime}}\mathbf{v}^{*}(s^{\prime})\right\}.$
Prove that $\pi$ is optimal.
3.4. Let Markov chain $P$ be reversible with respect to a distribution $\boldsymbol{\pi}$}, i.e.
@@
3.5. Suppose Value Iteration is initialized at an arbitrary vector $\mathbf{v}_{0}$ satisfying
$\mathbf{v}_{0}(s)\leq\max_{a\in A}\left\{R_{sa}+\gamma\sum_{s^{\prime}\in S}P^{a}_{ss^{\prime}}\mathbf{v}_{0}(s^{\prime})\right\}\qquad\text{for all }s\in S.$
Prove that the successive iterates $\mathbf{v}_{t}$ increase monotonically, that is, $\mathbf{v}_{t+1}\geq\mathbf{v}_{t}$ for all $t$, and that $\mathbf{v}_{t}\leq\mathbf{v}^{\star}$.
@@ 3.5. Suppose Value Iteration is initialized at an arbitrary vector $\mathbf{v}_{0}$ satisfying
In Chapter 2, we observed that even simple questions about general dynamical systems are intractable. This motivates restricting attention to a special class of dynamical systems for which we can design efficient algorithms with provable guarantees.
@@
If the matrices ${A}_{t}={A}$ and ${B}_{t}={B}$ are fixed, we say that the dynamical system is linear time-invariant (LTI). We have already seen examples of LTI systems in Chapter 2.
@@
A more general class of linear dynamical systems arises when the matrices ${A}_{t}$ and ${B}_{t}$ change over time. These are called linear time-varying (LTV) systems. The class of LTV systems can express more general dynamics via linearization, as we show next.
@@ 4.1 Approximating General Dynamics via LTV Systems
Consider deterministic dynamics of the form
$\mathbf{x}_{t+1}=f(\mathbf{x}_{t},\mathbf{u}_{t}),$
where $f$ is twice continuously differentiable. Fix a nominal state-control trajectory $(\bar{\mathbf{x}}_{t},\bar{\mathbf{u}}_{t})$ satisfying
$\bar{\mathbf{x}}_{t+1}=f(\bar{\mathbf{x}}_{t},\bar{\mathbf{u}}_{t}),$
and define the deviation variables
$\delta\mathbf{x}_{t}=\mathbf{x}_{t}-\bar{\mathbf{x}}_{t},\qquad\delta\mathbf{u}_{t}=\mathbf{u}_{t}-\bar{\mathbf{u}}_{t}.$
A first-order Taylor expansion around $(\bar{\mathbf{x}}_{t},\bar{\mathbf{u}}_{t})$ gives
$f(\mathbf{x}_{t},\mathbf{u}_{t})=f(\bar{\mathbf{x}}_{t},\bar{\mathbf{u}}_{t})+J_{\mathbf{x}}f(\bar{\mathbf{x}}_{t},\bar{\mathbf{u}}_{t})(\mathbf{x}_{t}-\bar{\mathbf{x}}_{t})+J_{\mathbf{u}}f(\bar{\mathbf{x}}_{t},\bar{\mathbf{u}}_{t})(\mathbf{u}_{t}-\bar{\mathbf{u}}_{t})+r_{t},$
where $J_{\mathbf{x}}f(\bar{\mathbf{x}}_{t},\bar{\mathbf{u}}_{t})$ and $J_{\mathbf{u}}f(\bar{\mathbf{x}}_{t},\bar{\mathbf{u}}_{t})$ are the Jacobians of $f$ with respect to $\mathbf{x}$ and $\mathbf{u}$, evaluated at $(\bar{\mathbf{x}}_{t},\bar{\mathbf{u}}_{t})$, and where the remainder satisfies
$\|r_{t}\|=O\!\left(\|\delta\mathbf{x}_{t}\|^{2}+\|\delta\mathbf{u}_{t}\|^{2}\right)$
locally around the nominal trajectory. Writing
$A_{t}\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}J_{\mathbf{x}}f(\bar{\mathbf{x}}_{t},\bar{\mathbf{u}}_{t}),\qquad B_{t}\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}J_{\mathbf{u}}f(\bar{\mathbf{x}}_{t},\bar{\mathbf{u}}_{t}),$
and subtracting the nominal dynamics yields the deviation dynamics
$\delta\mathbf{x}_{t+1}=A_{t}\delta\mathbf{x}_{t}+B_{t}\delta\mathbf{u}_{t}+r_{t}.$
Thus, near a nominal trajectory, a smooth nonlinear dynamical system is well approximated by a time-varying linear system in the deviation variables. If the original nonlinear system also has additive perturbations, then those perturbations simply enter additively in the linearized deviation dynamics as well.
This derivation explains why in practice the class of time-varying linear dynamical systems is very expressive, at least locally for smooth dynamical systems. In chapter 13 we see another reduction to LDS from nonlinear dynamics.
@@ 4.2 Stabilizability of Linear Systems
Verifying stability of a system, and more so stabilizability, was shown to be intractable in our discussion of Chapter 2. For linear dynamical systems, however, the picture is quite different: stability and stabilizability can be efficiently verified. Furthermore, the theoretical analysis of this fact is elegant.
@@
Recall that a system is stabilizable if, in the absence of perturbations, from every starting state there exists a control policy that drives the state to zero. Determining whether a linear time-invariant system is stabilizable admits a particularly simple characterization. Henceforth, we denote by $\rho({A})$ the spectral radius of a matrix A{A}, that is,
$\rho({A})=\max\{|\lambda_{1}|,\ldots,|\lambda_{d}|\},$
the maximum modulus of all (possibly complex) eigenvalues. The stabilizability of a linear dynamical system can then be characterized as follows.
@@
We prove the sufficiency of this condition and leave the necessity as Exercise 4.6. Both directions rely on the following characterization of the spectral radius.
@@
$\rho({A})<1$ if and only if $\lim_{t\to\infty}{A}^{t}=0$. Furthermore, if $\rho({A})>1$, then for any norm,
$\lim_{t\to\infty}\|{A}^{t}\|=\infty.$
The proof of this fact is left as Exercise 4.6. Based on it, we can now prove the sufficiency part of Theorem 4.1.
@@
Suppose that there exists a matrix K{K} such that $\rho({A}+{B}{K})<1$. Choose the linear state-feedback controller $\mathbf{u}_{t}={K}\mathbf{x}_{t}$. In the noiseless system, the closed-loop dynamics are
$\mathbf{x}_{t+1}=({A}+{B}{K})\mathbf{x}_{t},$
and therefore
$\mathbf{x}_{t}=({A}+{B}{K})^{t}\mathbf{x}_{0}.$
By the fact above, $(A+BK)^{t}\to 0$ as $t\to\infty$. Hence
$\lim_{t\to\infty}\|\mathbf{x}_{t}\|=\lim_{t\to\infty}\|({A}+{B}{K})^{t}\mathbf{x}_{0}\|=0.$
Thus the system is stabilizable. ∎
@@ 4.2.1 Efficient Verification of Stabilizability
Stabilizability admits a convenient convex characterization. We use the following standard form of the discrete-time Lyapunov theorem: a matrix $M$ satisfies $\rho(M)<1$ if and only if there exists a matrix $P\succ 0$ such that
$MPM^{\top}\prec P.$
By scaling $P$, we may additionally impose $P\succeq I$.
Applying this to the closed-loop matrix $A+BK$ gives the following linear matrix inequality.
@@
There exists a matrix $K$ such that $\rho(A+BK)<1$ if and only if there exist matrices $P\in\mathcal{S}^{d_{x}\times d_{x}}$ and $L\in{\mathbb{R}}^{d_{u}\times d_{x}}$ such that
Moreover, any feasible pair $(P,L)$ yields a stabilizing controller $K=LP^{-1}$.
@@
Suppose first that $\rho(A+BK)<1$. By the discrete-time Lyapunov theorem, there exists $P\succ 0$ such that
$(A+BK)P(A+BK)^{\top}\prec P.$
After scaling, we may assume $P\succeq I$. By the Schur complement, this is equivalent to
$\begin{bmatrix}P&((A+BK)P)^{\top}\\ (A+BK)P&P\end{bmatrix}\succ 0.$
Setting $L=KP$ gives the displayed condition.
Conversely, suppose $P$ and $L$ satisfy the displayed condition, and define $K=LP^{-1}$. Another application of the Schur complement yields
$(A+BK)P(A+BK)^{\top}\prec P.$
Hence $\rho(A+BK)<1$ by the discrete-time Lyapunov theorem. ∎
Thus stabilizability can be checked efficiently via semidefinite feasibility, and any feasible solution directly produces a stabilizing controller.